OpenStudy (anonymous):

The graph of y = -3x^2- 2x - 5 has a minimum or maximum value?

OpenStudy (studygurl14):

graph it.

OpenStudy (studygurl14):

https://www.desmos.com/calculator

OpenStudy (studygurl14):

This is also a parabola, indicated by the x^2, and downward curving, indicated by the -3, so the graph has a maximum.

OpenStudy (freckles):

If you have y=ax^2 and a is negative, you have a max. If you have y=ax^2 and a is positive, you have a min.

OpenStudy (freckles):

For example: Example 1: y=ax^2 and a is positive then you have something that looks like:  You have a min on that graph. Example 2: y=ax^2 and a is negative then you have something that looks like:  You have a max on this graph.